OpenStudy (anonymous):

A random number generator draws at random with replacement from the digits 0, 1, 2, 3, 4, 5, 6, 7, 8, 9. Find the chance that the digit 5 appears on more than 11% of the draws, if: PROBLEM 2A : 1.0 POINTS 100 draws are made unanswered You have used 0 of 2 submissions PROBLEM 2B : 1.0 POINTS 1000 draws are made

OpenStudy (anonymous):

can any one reply pls

OpenStudy (anonymous):

pls some one reply

OpenStudy (nubeer):

ok i can give a shot but not so sure.. P(5)= (1/10) for 100 draws P(5 appears in 11% draw) =100C11* [(1/10)^11]*[(9/10)^(100-11)]

OpenStudy (kropot72):

The probability that the digit 5 appears on more than 11% of the draws if 1000 draws are made can be found from the normal approximation to the binomial distribution. The mean = np = 1000 * 1/10 = 100 The standard distribution is given by \[\sigma=\sqrt{np(1-p)}=\sqrt{90}=9.4868\] The z-score for 110 draws is \[z=\frac{X-\mu}{\sigma}=\frac{110-100}{9.4868}=1.054\] Now you need to refer to a standard normal distribution table to find the cumulative probability corresponding to z = 1.054. This will give the cumulative probability that the digit 5 will appear up to 110 times in 1000 draws. To find probability that the digit 5 appears more than 110 times in 1000 subtract the probability just found from 1.0000.

OpenStudy (kropot72):

The same method can be used to find the probability that the digit 5 appears on more than 11% of the draws if 100 draws are made.

OpenStudy (kropot72):

@dinakar Do you understand the method?

OpenStudy (anonymous):

pls can u reply the exact ans .i dnt have any idea abt ths method .i dint go through ecxrcies

OpenStudy (kropot72):

On OS we try to help students to get solutions rather than give answers. However I will help you further with the 1000 draws question. A standard normal distribution table gives the cumulative probability of 0.8540 for a z-score of 1.054. The required probability is 1.0000 - 0.8540 = 0.146 or 14.6%

OpenStudy (anonymous):

ans 14.6% went wrong

OpenStudy (anonymous):

1)0.297 2)0.135

OpenStudy (kropot72):

The accuracy of the approximation can be improved by applying a 'continuity correction' of 0.5 to the calculation of the z-score. The z-score then becomes \[z=\frac{110+0.5-100}{9.4868}=1.107\] The required probability is 1.0000 - 0.8786 = 0.1214 or 12.14%

OpenStudy (anonymous):

thanks fr reply

OpenStudy (anonymous):

A die is rolled 12 times. Find the chance that the face with six spots appears once among the first 6 rolls, and once among the next 6 rolls.

OpenStudy (anonymous):

can any one post ans

OpenStudy (anonymous):

I toss a coin 4 times. Find the chance of getting 2heads

OpenStudy (anonymous):

actualy i have tried it ,i got 0.25 ans but it went wrong .pls any one can xplain

OpenStudy (anonymous):

ans;p(HH)=1/2*1/2=0.25.but it went wrong .pls help me out

OpenStudy (anonymous):

plssss someone help me out yar